OpenStudy (anonymous):

I need to prove this is a natural number, but don't say the answer please, I just want the first step.

OpenStudy (anonymous):

\[\sqrt{7 + 4\sqrt{3}} + \sqrt{7 - 4\sqrt{3}}\]

OpenStudy (anonymous):

Start by multiplying the expression by \[\sqrt{7-\sqrt{3}} \div \sqrt{7-\sqrt{3}}\]

OpenStudy (anonymous):

*sorry, the root 3's should be 4*root 3

OpenStudy (anonymous):

Okay, and all negative?

OpenStudy (anonymous):

yes, you're multiplying the expression by 1, but then stuff starts canceling out :)

OpenStudy (anonymous):

Can I call 7 a and 4qsrt3 b?

OpenStudy (anonymous):

If that makes it easier for you, I guess. So long as you realize a^2-b^2=1 :)

OpenStudy (anonymous):

Okay, let me try

OpenStudy (anonymous):

Omg, I got it! It's 14. Idk if it's right but it looks right

OpenStudy (anonymous):

Aww according to wolfram it is 4. At least I got close

OpenStudy (anonymous):

Can you help me, @SmoothMath ?

OpenStudy (anonymous):

I can show you my work if you want